OpenStudy (anonymous):

Can someone help me please. Rosa invests $3300 in an account with an APR of 3% and annual compounding. Julian invests $2800 in an account with an APR of 4% and annual compounding. A. After 5 years Rosa will have a balance of approximately $ ____. After 20 years Rosa will have a balance of approximately $_____. (Round to the nearest cent.)

OpenStudy (anonymous):

Question sounds familiar. Did you ask this last night? \[3300\times1.03^{5}\] = 3825.60 according to my calculator. Raise to the 20th power instead of the 5th power to answer the second part. I guess the stuff about Julian is irrelevant or perhaps there are later questions about him?

OpenStudy (anonymous):

For Rosas' investment. for 5 years. n=5, r= 3, P=$3300\[A=P(1+\frac{r}{100})^n = 3300(1+\frac{3}{100})^5 =3300(1+0.03)^5=3300(1.03)^5\]\[=3300(1.03)^5=3300 \times 1.159 =3824.7 - dollars\]

OpenStudy (anonymous):

So thats Rosas balance after 5 years?

OpenStudy (anonymous):

Yes....

OpenStudy (anonymous):

I disagree with the last line on dpasingh's reply. Do NOT round the 1.03^5 to 1.159. Don't round until your final answer.

OpenStudy (anonymous):

So 3,825.60

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

for 20 years do i just change the power from 5 to 20?

OpenStudy (anonymous):

You got it! I don't even think about the long involved formula for annual compounding. Just multiply by 1.0interest rate raised to however many years you wanna compound

OpenStudy (anonymous):

Can someone help me complete the rest of this problem. Rosa invests 3300 in an account with an APR of 3% and annual compounding. Julian invests 2800 in an account with an APR of 4% and annual compounding. Complete 1 through c. a. after 5 years Rosa will have a balance of approximately 3,825.60/ After 20 years Rosa will have a balance of approximately 5960.17. After 5 years Julian will have a balance of $____. After 20 years Julian will have a balance of approximately $ ___.